# Combinatorics and Hodge theory of degenerations of abelian varieties: A survey of the Mumford construction

### v2

@@ 6. The second Voronoi fan and Alexeev’s theorem
is integer-valued on ${\bf M}\simeq{\mathbb{Z}}^{g}$. For instance, we may take the coefficients of $L_{i}$ to be the diagonal entries of the matrix $B_{i}$ in some basis. Then, there is a unique convex section $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$ admitting a lift to $\bf M$$_{\mathbb{R}}$ which agrees with the above function $b_{i}$ on $\bf M$}.
It is a simple verification from the definition of the Delaunay cells that $\bigcup_{i}{\rm Bend}(\overline{b}_{i})$ is exactly ${\rm Del}_{\sum r_{i}B_{i}}$ for any $(r_{1},\dots,r_{k})\in{\mathbb{N}}^{k}$—this condition translates into a condition that the bending locus of the convex $\mathbb{Z}$}PL function which agrees with ${\bf m}\mapsto\sum_{i=1}^{k}r_{i}b_{i}({\bf m})$ on $\bf M$} is the same for all $(r_{1},\dots,r_{k})\in{\mathbb{N}}^{k}$. Thus, the $\overline{b}_{i}$ are dicing. Furthermore, the additional hypothesis of Construction 3.25 is satisfied: The associated bilinear forms $B_{i}$ define the extremal rays of a polyhedral cone in $({\mathcal{A}}_{g})_{\rm trop}$. Thus, we get an extension of the universal family
$X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$
which is proper over the deepest stratum of the toroidal boundary of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. $\clubsuit$
@@
The issue begins in the interior ${\mathcal{A}}_{g}$, see e.g. the discussion in [32, Sec. 1] and [49, Sec. 19 and bottom of p. 209]: For a given abelian variety $(A,0,L)$ with origin $0\in A$ and principal polarization $L\in{\rm NS}(A)$, there are $2^{2g}$ different possible (1)(-1)-symmetric lifts ${\mathcal{L}}\in{\rm Pic}(A)$ of $L$. These lifts define naturally a torsor over the $2$-torsion subgroup $A[2]$ and thus, on the universal family ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$ we have a natural torsor ${\rm Lifts}(L)\to{\mathcal{A}}_{g}$ under the group scheme ${\mathcal{X}}_{g}[2]\to{\mathcal{A}}_{g}$ of relative $2$-torsion in ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$. But ${\rm Lifts}(L)$ admits no section—it is impossible to globally lift $L$ to some (1)(-1)-symmetric ${\mathcal{L}}\in{\rm Pic}({\mathcal{X}}_{g}/{\mathcal{A}}_{g})$ when $g\geq 2$.
@@
For a cone ${\mathbb{B}}\in{\mathfrak{F}}_{\rm vor}$, Construction 6.5 gives a Mumford construction $X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. In the category of DM analytic stacks, this family descends as a family of polarized varieties over a tubular neighborhood of the deepest stratum of ${\mathcal{A}}_{g}^{{\mathbb{B}}}$. The reason is that $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$ which define the relevant polytopal Mumford Construction 3.25 are defined canonically by the $B_{i}$—one may worry that some non-canonicity is introduced by the choice of the characteristic linear form $L_{i}({\bf m})$ in (31). But a different choice $L_{i}\mapsto L_{i}+2L_{i}^{\prime}$ produces the same section $\overline{b}_{i}$.
On the other hand, due to the shift $L_{i}({\bf m})$, the resulting Mumford construction has no canonical origin section. Thus, the output of Construction 6.5 does not glue canonically to the universal family ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$ (which has an origin section), but rather to the universal family ${\mathcal{X}}_{g}^{\star}\to{\mathcal{A}}_{g}$ (which has a canonical lift of the principal polarization). If one were to take $L_{i}=0$ in (31), we would retain a canonical origin point, but the $\overline{b}_{i}$ would only be $\tfrac{1}{2}$-dicing, see [49].
By their canonicity and the uniqueness of gluings, these Mumford constructions are compatible between adjacencies of cones in ${\mathfrak{F}}_{\rm vor}$. Thus, we may glue them to produce an extension
$\overline{{\mathcal{X}}_{g}^{\star}}^{\rm vor}\coloneqq{\mathcal{X}}_{g}^{\star}\cup\textstyle\bigcup_{\,{\mathbb{B}}\,\in\,\operatorname{GL}_{g}({\mathbb{Z}})\backslash{\mathfrak{F}}_{\rm vor}}X^{\rm univ}(\mathbbm{b})\to\overline{{\mathcal{A}}}_{g}^{\rm vor}$
which is proper over a neighborhood of the deepest toroidal stratum of the Baily–Borel boundary. In fact, the construction extends even over the Baily–Borel strata of intermediate dimension; we refer the interested reader to [49, Sec. 13, Prop. 13.5]. Roughly, this follows from the fact that the analytic tubular neighborhood $T(\mathbbm{b})\subset Y(\mathbbm{b})$ of the deepest stratum in Construction 3.25 may be chosen to surject onto the uniformization ${\mathcal{H}}_{h}\to{\mathcal{A}}_{h}$ of the Baily–Borel cusp, and we may then further quotient the extended universal family by ${\rm Sp}_{2h}({\mathbb{Z}})$.
@@
extending the universal family of abelian varieties. But it follows from Serre’s GAGA for Deligne–Mumford stacks, see [60, Cor. 5.13], that $\overline{{\mathcal{X}}_{g}^{\star}}^{\rm vor}$ is a DM algebraic stack and $f_{\rm vor}$ is projective. For instance, $f_{\rm vor}$ becomes a morphism of projective schemes after taking the pullback toroidal compactification of an appropriate étale cover.
@@
of the universal pair $(X,\Theta)$. Here, the theta divisor $\Theta\in|{\mathcal{L}}|$ is the unique element of the linear system. Note that $\Theta$ extends as an effective, relatively ample divisor over $\overline{{\mathcal{A}}}_{g}^{\rm vor}$ as the vanishing locus of the unique weight $w=1$ theta function $\Theta=V(\Theta_{(0/1,\dots,0/1)})$ of Construction 3.20. $\clubsuit$
@@
There is one orbit each of $3$-, 22-, 11-, and 0-dimensional cones, corresponding respectively to the cographic cones of the theta graph, the wedge of two circles, a single circle, and a point. Over a maximal, $3$-dimensional cone, the universal family $\overline{{\mathcal{X}}^{\star}_{2}}^{\rm vor}\to\overline{{\mathcal{A}}}_{2}^{\rm vor}$ is extended by a Mumford construction $X(\mathbbm{b})\to\Delta^{3}$ isomorphic to Example 3.34. The reduction theory, i.e. analysis of $\operatorname{GL}_{2}({\mathbb{Z}})$-equivalence classes, of positive-definite bilinear forms of rank $2$, goes back at least to work of Fricke–Klein [37]; see Vallentin [62, Ch. 2] for some historical discussion.
@@
$g=4$: Here, the second Voronoi fan is ${\mathfrak{F}}_{\rm vor}=\operatorname{GL}_{4}({\mathbb{Z}})\cdot\{{\mathbb{B}}_{\rm black},\,{\mathbb{B}}_{\rm grey},\,{\mathbb{B}}_{\rm white}\}$, see for instance [62, Sec. 4.4.1]. That is, there are three $\operatorname{GL}_{4}({\mathbb{Z}})$-orbits of maximal, $10$-dimensional cones of ${\mathfrak{F}}_{\rm vor}$. The original computation goes back to Delaunay [24, Thm. III], who found all but one of the $\operatorname{GL}_{4}({\mathbb{Z}})$-orbits of cones of ${\mathfrak{F}}_{\rm vor}$, and Shtogrin [58], who filled the gap.
@@
The number of orbits of cones of dimensions $10$, 99, 88, 77, 66, 55, 44, 33, 22, 11, 0 are, respectively, $3$, 44, 77, $11$, $11$, 99, 77, 44, 22, 22, 11, see Hulek–Tommasi [35, p. 232]. In particular, the toroidal compactification $\overline{{\mathcal{A}}}_{4}^{\rm vor}$ has $2$ boundary divisors and $3$ zero-dimensional strata.
@@
is an isomorphism. To see this, note first that by [54, Prop. XIII.2.3], the group $H^{1}({\mathcal{A}}_{g},{\mathcal{X}}_{g})$ is torsion. Second, for each $n\in{\mathbb{N}}$, the natural map $H^{1}({\mathcal{A}}_{g},{\mathcal{X}}_{g}[n])\to H^{1}({\mathcal{A}}_{g},{\mathcal{X}}_{g})[n]$ is an isomorphism by the long exact sequence in cohomology arising from the short exact sequence
@@
Let $\epsilon$ be a sufficiently small positive rational number. Then every fiber of $(\overline{{\mathcal{X}}_{g}^{\star}}^{\rm vor},\epsilon\overline{\Theta}_{g}^{\rm vor})\to\overline{{\mathcal{A}}}_{g}^{\rm vor}$ is a KSBA-stable pair.
@@
In any Mumford construction of second Voronoi type, see Construction 6.5, all fibers have slc singularities, and the canonical bundle $K_{X}\simeq{\mathcal{O}}_{X}$ is trivial. This follows from a mild generalization to affine toric bases of Proposition 3.13, by checking that slices ${\mathcal{S}}_{(r_{1},\dots,r_{k})}$ of the normal fan, for $r_{1}B_{1}+\cdots+r_{k}B_{k}$ integral, are integral tilings of $\bf N$$_{\mathbb{R}}$. Indeed, ${\mathcal{S}}_{(r_{1},\dots,r_{k})}$ is, up to translation, the image of the Voronoi decomposition ${\rm Vor}_{B}$ under the map $N_{\mathbb{R}}\colon{\bf M}_{\mathbb{R}}\to{\bf N}_{\mathbb{R}}$ corresponding to $B=\sum_{i=1}^{k}r_{i}B_{i}$. A linear algebra computation verifies the integrality.
The log canonical centers of $X$ are exactly the toric strata of the Mumford construction. Given an effective divisor $D\subset X$, there is an $\epsilon\ll 1$ for which $(X,\epsilon D)$ defines a KSBA-stable pair if and only if $D$ contains no log canonical centers, i.e. toric strata. In fact, in our setting, any $\epsilon\leq 1$ suffices, see [3, Thm. 3.10], generalizing [38, Thm. 17.13].
We claim this property follows from the definition of $\Theta_{(0/1,\,\dots,\,0/1)}$. The key observation is that, for every vertex of a polyhedral face $F\subset{\rm Del}_{B}$, $B\in{\mathbb{B}}^{\circ}$, the restriction $\Theta_{(0/1,\,\dots,\,0/1)}|_{Y_{F}}$ of the theta divisor to the stratum $Y_{F}\subset X$ is a section of a (toric) line bundle, for which the coefficient of any monomial corresponding a vertex of $F$ is nonzero, i.e. lies in $\mathbb{C}$$^{*}$. This property ensures that the restriction of the theta divisor contains no toric strata of $Y_{F}$. ∎
@@
For $\epsilon\ll 1$, $(\overline{{\mathcal{X}}_{g}^{\star}}^{\rm vor},\epsilon\overline{\Theta}_{g}^{\rm vor})\to\overline{{\mathcal{A}}}_{g}^{\rm vor}$ is the universal family over the normalization of the KSBA compactification of the space of KSBA-stable pairs $(X,\epsilon\Theta)$, with $X$ a torsor under a $g$-dimensional PPAV and $\Theta\subset X$ the theta divisor.
@@
By Proposition 6.10, there is a classifying morphism $c\colon\overline{{\mathcal{A}}}_{g}^{\rm vor}\to\overline{{\mathcal{A}}}_{g}^{\Theta}$ where the latter is, by definition, the closure, taken with reduced scheme structure, of the space of pairs $(X,\epsilon\Theta)$ as in Construction 6.6, in the separated DM stack of KSBA-stable pairs [39, 3, 40]. By Zariski’s main theorem and the normality of toroidal compactifications, it suffices to check that $c$ is finite.

### v3

@@ 6.1. The universal family of abelian torsors with theta divisor
is integer-valued on ${\bf M}\simeq{\mathbb{Z}}^{g}$. For instance, we may take the coefficients of $L_{i}$ to be the diagonal entries of the matrix $B_{i}$ in some basis. Then, there is a convex section $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Q}}{\rm PL}/{\mathbb{Q}}{\rm L})$ admitting a lift to $\bf M$$_{\mathbb{R}}$ which agrees with the above function $b_{i}$ on $\bf M$} (in contrast to the sheaf $\tfrac{1}{d}{\mathbb{Z}}{\rm PL}$, sections of $\mathbb{Q}$}PL allow all $a_{i}\in{\mathbb{Q}}$ in the notation of Definition 3.25, see (18)).
Since $b_{i}({\bf m})$ is integer-valued on $\bf M$}, the function $b_{i}$ is ${\mathbb{Z}}{\rm PL}$ on each cell of the Delaunay decomposition that generates $\bf M$}, in the sense that the differences between all vertices of the cell generate $\bf M$}. When $g\leq 4$, every Delaunay cell is generating, but once $g\geq 5$, there are counterexamples, see e.g. [28, p. 796] and [7, Sec. 1.14, Ex. 1.15].
It is a simple verification from Definition 6.1 that $\bigcup_{i}{\rm Bend}(\overline{b}_{i})$ is exactly ${\rm Del}_{\sum r_{i}B_{i}}$ for any $(r_{1},\dots,r_{k})\in{\mathbb{N}}^{k}$—this condition translates into a condition that the bending locus of the convex $\mathbb{Q}$}PL function which agrees with ${\bf m}\mapsto\sum_{i=1}^{k}r_{i}b_{i}({\bf m})$ on $\bf M$} is the same for all $(r_{1},\dots,r_{k})\in{\mathbb{N}}^{k}$. Thus, the $\overline{b}_{i}$ are dicing. Furthermore, the additional condition of Construction 3.26 is satisfied: The associated bilinear forms $B_{i}$ span extremal rays of a polyhedral cone in $({\mathcal{A}}_{g})_{\rm trop}$. Despite $b_{i}$ only being ${\mathbb{Q}}{\rm PL}$, we may still take the overgraph $\Gamma$ of the $b_{i}$ as a polytope, and form its normal fan, see Constructions 3.11, 3.26. Thus, we get a relatively proper extension of the universal family $X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}.$ Note that not all fibers need be reduced, see Remark 3.4. $\clubsuit$
@@
The issue begins in the interior ${\mathcal{A}}_{g}$, see e.g. the discussion in [33, Sec. 1] and [51, Sec. 19 and bottom of p. 209]: For a given abelian variety $(A,0,L)$ with origin $0\in A$ and principal polarization $L\in{\rm NS}(A)$, there are $2^{2g}$ different possible (1)(-1)-symmetric lifts ${\mathcal{L}}\in{\rm Pic}(A)$ of $L$. These lifts define naturally a torsor over the $2$-torsion subgroup $A[2]$ and thus, on the universal family ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$ we have a natural torsor ${\rm Lifts}(L)\to{\mathcal{A}}_{g}$ under the group scheme ${\mathcal{X}}_{g}[2]\to{\mathcal{A}}_{g}$ of relative $2$-torsion in ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$. But ${\rm Lifts}(L)$ admits no section—it is impossible to globally lift $L$ to some (1)(-1)-symmetric ${\mathcal{L}}\in{\rm Pic}({\mathcal{X}}_{g}/{\mathcal{A}}_{g})$ when $g\geq 2$.
@@
For a cone ${\mathbb{B}}\in{\mathfrak{F}}_{\rm vor}$, Construction 6.5 gives a Mumford construction $X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. In the category of DM analytic stacks, this family descends as a family of polarized varieties over an étale neighborhood of the boundary strata of ${\mathcal{A}}_{g}^{{\mathbb{B}}}$. The reason is that $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Q}}{\rm PL}/{\mathbb{Q}}{\rm L})$ which define the relevant polytopal Mumford Construction 3.26 are defined canonically by the $B_{i}$—one may worry that some non-canonicity is introduced by the choice of the characteristic linear form $L_{i}({\bf m})$ in (36) which determine the lifts $b_{i}$. But a different choice $L_{i}\mapsto L_{i}+2L_{i}^{\prime}$ produces the same section $\overline{b}_{i}$.
On the other hand, due to the shifts $L_{i}({\bf m})$, the resulting Mumford construction has no canonical origin section, see Remark 3.39. Thus, the output of Construction 6.5 does not glue canonically (i.e. in a manner independent of the choice of $L_{i}$) to the universal family ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$ (which has an origin section), but rather to the universal family ${\mathcal{X}}_{g}^{\star}\to{\mathcal{A}}_{g}$ (which has a canonical lift of the principal polarization). If one were to take $L_{i}=0$ in (36), we would retain a canonical origin point, but the principal polarization may not lift to a theta divisor, see [51].
By their canonicity and the uniqueness of gluings, the Mumford constructions of Construction 6.5 are compatible between adjacencies of cones in ${\mathfrak{F}}_{\rm vor}$. Thus, we may glue them via the unique gluings respecting the lift of $L$, to produce a proper extension
$\overline{{\mathcal{X}}_{g}^{\star}}^{\rm vor}\coloneqq{\mathcal{X}}_{g}^{\star}\cup\textstyle\bigcup_{\,{\mathbb{B}}\,\in\,\operatorname{GL}_{g}({\mathbb{Z}})\backslash{\mathfrak{F}}_{\rm vor}}X^{\rm univ}(\mathbbm{b})\to\overline{{\mathcal{A}}}_{g}^{\rm vor}.$
@@
extending the universal family of abelian torsors with lift of principal polarization. It follows from Serre’s GAGA for Deligne–Mumford stacks, see [64, Cor. 5.13], that $\overline{{\mathcal{X}}_{g}^{\star}}^{\rm vor}$ is a DM algebraic stack and $f_{\rm vor}$ is projective. For instance, $f_{\rm vor}$ becomes a morphism of projective schemes after taking the pullback toroidal compactification of an appropriate étale cover.
@@
of the universal pair $(X,\Theta)$. Here, the theta divisor $\Theta\in|{\mathcal{L}}|$ is the unique element of the linear system. Note that $\Theta$ extends as an effective, relatively ample divisor over $\overline{{\mathcal{A}}}_{g}^{\rm vor}$ as the vanishing locus of the unique weight $w=1$ theta function $\Theta=V(\Theta_{(0/1,\dots,0/1)})$ of Construction 3.21. $\clubsuit$
@@
There is one orbit each of $3$-, 22-, 11-, and 0-dimensional cones, corresponding respectively to the cographic cones of the theta graph, the wedge of two circles, a single circle, and a point. Over a maximal, $3$-dimensional cone, the universal family $\overline{{\mathcal{X}}^{\star}_{2}}^{\rm vor}\to\overline{{\mathcal{A}}}_{2}^{\rm vor}$ is extended by a Mumford construction $X(\mathbbm{b})\to\Delta^{3}$ isomorphic to Example 3.35. The reduction theory, i.e. analysis of $\operatorname{GL}_{2}({\mathbb{Z}})$-equivalence classes, of positive-definite bilinear forms of rank $2$, goes back at least to work of Fricke–Klein [38]; see Vallentin [67, Ch. 2] for some historical discussion.
@@
$g=4$:Here, the second Voronoi fan is ${\mathfrak{F}}_{\rm vor}=\operatorname{GL}_{4}({\mathbb{Z}})\cdot\{{\mathbb{B}}_{\rm black},\,{\mathbb{B}}_{\rm grey},\,{\mathbb{B}}_{\rm white}\}$, see for instance [67, Sec. 4.4.1]. That is, there are three $\operatorname{GL}_{4}({\mathbb{Z}})$-orbits of maximal, $10$-dimensional cones of ${\mathfrak{F}}_{\rm vor}$. The original computation goes back to Delaunay [24, Thm. III], who found all but one of the $\operatorname{GL}_{4}({\mathbb{Z}})$-orbits of cones of ${\mathfrak{F}}_{\rm vor}$, and Shtogrin [61], who filled the gap.
@@
The number of orbits of cones of dimensions $10$, 99, 88, 77, 66, 55, 44, 33, 22, 11, 0 are, respectively, $3$, 44, 77, $11$, $11$, 99, 77, 44, 22, 22, 11, see Hulek–Tommasi [36, p. 232]. In particular, the toroidal compactification $\overline{{\mathcal{A}}}_{4}^{\rm vor}$ has $2$ boundary divisors and $3$ zero-dimensional strata.
@@
is an isomorphism. To see this, note first that by [57, Prop. XIII.2.3], the group $H^{1}({\mathcal{A}}_{g},{\mathcal{X}}_{g})$ is torsion. Second, for each $n\in{\mathbb{N}}$, the natural map $H^{1}({\mathcal{A}}_{g},{\mathcal{X}}_{g}[n])\to H^{1}({\mathcal{A}}_{g},{\mathcal{X}}_{g})[n]$ is an isomorphism by the long exact sequence in cohomology arising from the short exact sequence
@@
Let $\epsilon>0$ be a sufficiently small positive rational number. There is a canonical stacky toroidal compactification ${\mathcal{A}}_{g}\hookrightarrow\overline{{\mathcal{A}}}_{g}^{\rm nvor}$ admitting a partial coarsening $\overline{{\mathcal{A}}}_{g}^{\rm nvor}\to\overline{{\mathcal{A}}}_{g}^{\rm vor}$ along which every fiber of the normalized pullback $(\overline{{\mathcal{X}}_{g}^{\star}}^{\rm nvor},\epsilon\overline{\Theta}_{g}^{\rm nvor})\to\overline{{\mathcal{A}}}_{g}^{\rm nvor}$ of Construction 6.6 is a KSBA-stable pair.
The toric stack structure corresponds to “toric stacky data” in the sense of [66, Def. 4.1]: For the cones ${\mathbb{B}}\in{\mathfrak{F}}_{\rm vor}$ we require the additional data of $\operatorname{GL}_{g}({\mathbb{Z}})$-equivariant, finite index sublattices of ${\mathbb{R}}{\mathbb{B}}\cap{\rm Sym}_{g\times g}({\mathbb{Z}})$, which are compatible along face inclusions of cones in ${\mathfrak{F}}_{\rm vor}$.
@@
It follows from a mild generalization of Proposition 3.14 to affine toric bases that in a Mumford fan Construction 3.11, the fibers are reduced with trivial canonical bundle $K_{X}\simeq{\mathcal{O}}_{X}$ and slc singularities, whenever the slices ${\mathcal{S}}_{(r_{1},\dots,r_{k})}$ of the normal fan, for $r_{1}B_{1}+\cdots+r_{k}B_{k}\in{\rm Sym}_{g\times g}({\mathbb{Z}})$ integral, are integral tilings of $\bf N$$_{\mathbb{R}}$. For the Mumford Construction 6.5 of second Voronoi type, ${\mathcal{S}}_{(r_{1},\dots,r_{k})}$ is, up to translation, the image of the Voronoi decomposition ${\rm Vor}_{B}$ under the map $N_{\mathbb{R}}\colon{\bf M}_{\mathbb{R}}\to{\bf N}_{\mathbb{R}}$ corresponding to $B=\sum_{i=1}^{k}r_{i}B_{i}$.
In general, the desired integrality condition fails for $g\geq 5$, see [7, Sec. 1.14], exactly because the $b_{i}$ of Construction 6.6 may only be ${\mathbb{Q}}{\rm PL}$. But we may define canonically a finite index sublattice of ${\mathbb{R}}{\mathbb{B}}\cap{\rm Sym}_{g\times g}({\mathbb{Z}})$ as the one generated by the linear combinations $r_{1}B_{1}+\cdots+r_{k}B_{k}$ for which the slices ${\mathcal{S}}_{(r_{1},\dots,r_{k})}$ are integral polytopes. This defines toric stacky data, which is $\operatorname{GL}_{g}({\mathbb{Z}})$-equivariant, and so determines a DM stack $\overline{{\mathcal{A}}}_{g}^{\rm nvor}$ admitting $\overline{{\mathcal{A}}}_{g}^{\rm vor}$ as a partial coarsening. See also [44]. The normalized pullback of the Mumford Construction 6.5 along the partial coarsening map $\overline{{\mathcal{A}}}_{g}^{\rm nvor}\to\overline{{\mathcal{A}}}_{g}^{\rm vor}$ is now flat with reduced fibers.
The log canonical centers of a fiber $X^{n}$ of this pullback are exactly the toric strata of the Mumford construction. Given an effective divisor $D\subset X^{n}$, there is an $\epsilon\ll 1$ for which $(X^{n},\epsilon D)$ defines a KSBA-stable pair if and only if $D$ contains no log canonical centers, i.e. toric strata. In fact, in our setting, any $\epsilon\leq 1$ suffices, see [3, Thm. 3.10], generalizing [39, Thm. 17.13].
We claim this property follows from the definition of $\Theta_{(0/1,\,\dots,\,0/1)}$. The key observation is that, for every vertex of a polyhedral face $F\subset{\rm Del}_{B}$, $B\in{\mathbb{B}}^{\circ}$, the restriction $\Theta_{(0/1,\,\dots,\,0/1)}|_{Y_{F}}$ of the theta divisor to the stratum $Y_{F}\subset X^{n}$ is a section of a (toric) line bundle, for which the coefficient of any monomial corresponding a vertex of $F$ is nonzero, i.e. lies in $\mathbb{C}$$^{*}$. This property ensures that the restriction of the theta divisor contains no toric strata of $Y_{F}$. ∎
We remark that the procedure of choosing a finite index subcone of ${\mathbb{B}}\cap{\rm Sym}_{g\times g}({\mathbb{Z}})$ is unnecessary for matroidal cones, as in this setting, the Voronoi cell ${\mathcal{S}}_{(r_{1},\dots,r_{k})}$ is always integral—it is a Minkowski sum of integral line segments $0,r_{i}{\bf x}_{i}]$.
@@
For $\epsilon\ll 1$, $(\overline{{\mathcal{X}}_{g}^{\star}}^{\rm nvor},\epsilon\overline{\Theta}_{g}^{\rm nvor})\to\overline{{\mathcal{A}}}_{g}^{\rm nvor}$ is the universal family over the normalization of the KSBA compactification of the space of KSBA-stable pairs $(X,\epsilon\Theta)$, with $X$ a torsor under a $g$-dimensional PPAV and $\Theta\subset X$ the theta divisor.
@@
We will only address the equality of coarse spaces.
By Proposition 6.10, there is a classifying morphism $c\colon{\overline{A}}_{g}^{\rm vor}\to{\overline{A}}_{g}^{\Theta}$ where the latter is, by definition, the closure, taken with reduced scheme structure, of the coarse space of pairs $(X,\epsilon\Theta)$ as in Construction 6.6 in the moduli space of KSBA-stable pairs [40, 3, 41]. By Zariski’s main theorem and the normality of toroidal compactifications, it suffices to check that $c$ is finite.
@@
6.2. Algebraicity and projectivity
We now analyze under what circumstances an extension of the universal family ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$ or ${\mathcal{X}}_{g}^{\star}\to{\mathcal{A}}_{g}$ of principally polarized abelian varieties or torsors, by Mumford constructions, are either algebraic or projective.